OpenStudy (anonymous):

y = tanh^-1 (1-x/1+x) get dy/dx

sam (.sam.):

y=tanh^(-1)(1-(x)/(1)+x) y=tanh^(-1)(1) (1)/(u^(2)+1) d/dx tanh^(-1)(1)=0 =0

OpenStudy (vishal_kothari):

@sam u r right....